A beam of light from a laser hits a grid with the grid constant of 1.67 microns. Behind the grid, 5 rays can be observed. The angle between the outermost rays was measured to 94.1 degrees. Calculate the wavelength of the laser light.
1
Expert's answer
2020-04-22T10:02:36-0400

n = 5

angle("\theta") = 94.1o

d = 1.67 microns = 1.67 × 10−6 m

dsin("\theta") = n × "\lambda"

λ = dsin("\theta") / n = (1.67 × 10−6 × sin(94.1)) / 5 = 0.33 × 10-6 m

Answer: λ = 0.33 × 10−6 m.